The balloon contains 0.30 mol of helium starts at a pressure of 110 kPa and rises to an altitude where the pressure is 55.0 kPa , maintaining a constant 300 K temperature.

A) By what factor does its volume increase?

B)How much work does the gas in the balloon do?
Answer #1

IDEAL GAS EQUATION

P1V1 = nRT

V1 = nRT/P1 = 0.3 *8.314*300/110*103 = 0.0068 m3

PART A)

Boyle's law states that at constant temperature for a fixed mass, the absolute pressure and the volume of a gas are inversely proportional. The law can also be stated in a slightly different manner, that the product of absolute pressure and volume is always constant.

PV = CONSTANT

P1V1 =P2V2

V2 = P1V1/P2 = 110*103*0.0068/55*103 = 0.0136 m3

PART B) WORK DONE

W = 0.30*8.315*300*ln(V2/V1) = 0.30*8.315*300*ln(P1/P2)
= 0.30*8.314*300*ln(115/55) = 518.65 J